OpenStudy (anonymous):

3 inches of rope equals 7 cents. How much is 2 feet?

OpenStudy (anonymous):

how many inches in a foot?

OpenStudy (igreen):

@Mcmichaeni

OpenStudy (anonymous):

Well there are 12 inches in a foot. Divide 12 by 3 which gives you four. Four times 3 is 12. So 7 times 12 =84 so three feet of rope is 84cents

OpenStudy (anonymous):

12/3=4, 7*4=28, 28*2=56 your answer is 56

OpenStudy (anonymous):

this answer is 84